# Causal Feedback Discovery using Convergence Cross Mapping on Sea Ice Data

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; November 03–06, 2025; Minneapolis, MN, USAisbn: 978-1-4503-XXXX-X/2025/12ccs: Applied computing Earth and atmospheric sciences
@@ 1. Introduction
The areal extent, concentration, and thickness of sea ice in the Arctic ocean and adjacent seas have dramatically decreased over recent decades (Stroeve et al., 2012). This decline has significantly increased heat flux from the ocean to the atmosphere in autumn and early winter, leading to locally increased air temperatures, moisture, and cloud cover, while reducing static stability in the lower troposphere. Studies suggest that with advancing global warming, cold winters in mid-latitude continents may become uncommon in the second half of the twenty-first century. Additionally, recent findings have linked sea ice decline to summer precipitation patterns in Europe, the Mediterranean, and East Asia (Vihma, 2014; Ali et al., 2022). The reduction of Arctic ice sheets is a prominent indicator of warming, evidenced by a nearly 50% decrease in ice extent since 1979 (Stammerjohn et al., 2012). As the Arctic sea ice cover diminishes, the annual mean 2-m air temperature has increased across nearly all weather stations north of 60N, particularly in coastal and archipelago areas surrounding the Arctic Ocean (Polyakov et al., 2012).
@@
The effects of sea ice decline can be categorized into local and remote effects depending on the season. Local effects occur in regions that have lost sea ice cover while remote effects impact areas that have historically been ice-free. The expected direct local effects include increased turbulent fluxes of sensible and latent heat from the ocean to the atmosphere, as well as enhanced longwave radiation emitted by the sea surface. Understanding the remote effects is more complex, as it involves discerning how mid-latitude circulation responds to Arctic changes in the absence of other influencing factors. Interactions among atmospheric time series variables are inherently complex and nonlinear due to the intricate and often unpredictable dependencies that exist between them. In the atmosphere, factors such as temperature, humidity, sea ice extent, and radiation do not evolve in isolation; rather, they interact through multifaceted pathways. For example, an increase in temperature can accelerate sea ice melt, which reduces surface albedo and subsequently amplifies local and regional warming, thereby altering weather patterns. The abundance of feedback loops and deterministic components in atmospheric processes causes information about these variables to become inseparably intertwined. Identifying causal relationships is therefore essential to disentangle these interactions, improve predictive understanding, and design effective interventions. These interactions are fundamentally nonlinear because small perturbations in one variable can trigger disproportionately large responses in others, rendering the system highly sensitive to initial conditions. Moreover, feedback mechanisms where the output of a process feeds back into its own input further amplify complexity and nonlinearity. In this context, uncovering causal relationships and feedback loops is critical for advancing climate science, enabling more accurate forecasts, and informing targeted mitigation and adaptation strategies. Recent work applying Convergent Cross Mapping (CCM) to Arctic datasets has demonstrated its ability to uncover nonlinear feedbacks between sea ice, atmospheric moisture, and radiation processes, validating its use in polar climate research (Su and Wong, 2023; Gao et al., 2023). While controlled randomized experiments are ideal for efficient causal discovery, they are often unethical, costly, or technically unfeasible (Hoyer et al., 2008). As a result, research has focused on causal discovery methods that infer relationships from uncontrolled data. Although stochastic causal models such as Granger causality, PCMCI and VarLiNGAM are frequently employed due to their simplicity, they often fail to accurately capture complex, non-linear causal relationships because they are fundamentally constrained by their assumptions of linearity, stationarity, and acyclicity whereas atmospheric and climate dynamics are nonlinear(feedbacks, thresholds, bifurcations), nonstationary (climate change alters statistical properties over time), cyclic and feedback-driven (e.g., ice–albedo, ocean–atmosphere coupling) and finally high-dimensional with hidden confounders.In nonlinear systems every coupled variable carry information about others. This implies that variables cannot be fully removed from the system for analysis, and this further violates the linearity assumption made by stochastic causal models.
@@ 2.1. Background
Detecting causal drivers and regime shifts in Arctic climate systems presents a critical challenge due to nonlinear feedbacks, high-dimensional couplings, and non-separable spatiotemporal dynamics. Sea ice extent, atmospheric circulation, and oceanic heat fluxes interact through bidirectional feedbacks, complicating conventional stochastic causal discovery methods. Convergent Cross Mapping (CCM) addresses these limitations by reconstructing state-space manifolds from observed time series and assessing whether the historical states of one variable can reliably estimate another, enabling detection of weak and asymmetric causal links without relying on predictive models. Concurrently, advances in deep unsupervised spatiotemporal representation learning (Nji et al., 2024b, a; Faruque et al., 2023) demonstrate the ability to uncover latent temporal regimes, structural shifts, and evolving patterns in high-dimensional climate data. While representation learning identifies emergent regime structures, CCM provides mechanistic insight by quantifying the directional couplings that drive these patterns, revealing feedback loops and causal pathways underlying Arctic amplification.
@@
CCM builds on Takens’ theorem (Takens, 2006) and the Empirical Dynamic Modeling (EDM) framework, which reconstructs system attractors from time-delayed coordinates, forming “shadow manifolds” that preserve the topological properties of the underlying system. Cross mapping exploits the fact that if variable $X$ causally influences $Y$, then information about $X$ is embedded in $Y$’s manifold, enabling $X$ to be predicted from $Y$. Unlike Granger causality, which tests predictive improvement, CCM explicitly detects state-space dependence, making it particularly robust in nonlinear, noisy systems. Applications in climate science confirm that CCM successfully identifies causal feedbacks between sea ice extent, temperature, and radiative processes (Su and Wong, 2023; Gao et al., 2023), offering a powerful framework for disentangling drivers of Arctic sea ice decline and associated regime shifts.
The CCM Algorithm
Let two time series X and Y be defined as:
$X=\{X(1),X(2),\dots,X(L)\},\quad Y=\{Y(1),Y(2),\dots,Y(L)\},$
where $L$ is the length of the time series. The CCM procedure is as follows:
  1. (1)
    Reconstruct lagged-coordinate vectors. For embedding dimension $E$ and lag $\tau$, construct $\bar{x}(t)=\langle X(t),X(t-\tau),X(t-2\tau),\dots,X(t-(E-1)\tau)\rangle,$ for $t\in[1+(E-1)\tau,L]$.
  2. (2)
    Form the shadow manifold. Define the attractor manifold $M_{X}=\{\bar{x}(t)\,|\,t\in[1+(E-1)\tau,L]\}.$
  3. (3)
    Nearest neighbors. For each $\bar{x}(t)\in M_{X}$, find the $E+1$ nearest neighbors $\{\bar{x}(t_{1}),\dots,\bar{x}(t_{E+1})\}$ in Euclidean space. The $E+1$ requirement ensures a minimal simplex in $E$ dimensions.
  4. (4)
[…]
  5. (5)
    Predict $Y$ from $M_{X}$. Construct the CCM estimator: $\hat{Y}|M_{X}=\sum_{i=1}^{E+1}w_{i}\,Y(t_{i}).$
  6. (6)
    Test for convergence. If $X$ and $Y$ are dynamically coupled, then neighborhoods in $M_{X}$ correspond to neighborhoods in $M_{Y}$. As $L\to\infty$, $\hat{Y}|M_{X}$ should converge to $Y$.
  7. (7)
    Causal inference. Compute the correlation between $Y$ and $\hat{Y}|M_{X}$. Significant correlation implies that $Y$ influences $X$, since information about $Y$ is embedded in $X$. Symmetrically, repeating the procedure in the opposite direction tests if $X$ influences $Y$.
Thus, CCM establishes causality by detecting whether cross-mapped estimates converge with increasing data length, distinguishing true causal interactions from spurious correlations.
@@ 2.2. Related Work
Convergent Cross Mapping is designed to identify and quantify causalities in systems whose variables are not separable. Clark et al. proposed Multispatial CCM which combine the existing techniques of CCM and dewdrop regression to build a novel test of causal relations that leverages spatial replication to detect causal relationships between processes (Clark et al., 2015). While the ability of their proposed model was tested on simulated and real-world ecological data, their performance on real world spatiotemporal data is yet unexplored. Brouwer et al. (De Brouwer et al., 2020), proposed Latent CCM which uses reconstruction between latent processes of dynamical systems to infer causality between short and sporadic time series. Although their proposed seemed to work well on neural activity data, its application to climate data has not yet been explored. Ye et. al. (Ye et al., 2015), demonstrate the ability of CCM to identify different time-delayed interactions, distinguish between synchrony induced by strong unidirectional-forcing and true bidirectional causality, and resolve transitive causal chains.
@@ 3. Methodology
Our goal is to detect both linear and non-linear causal relationships, including weak and bidirectional feedback loops, among time series atmospheric variables. To achieve this, we evaluate a set of causal discovery methods without bias: stochastic models (Granger causality (Barnett et al., 2009), PCMCI (Runge, 2020), VarLiNGAM (Shimizu et al., 2006)) and the nonlinear Convergent Cross Mapping (CCM), applied to both synthetic and real-world Arctic climate data.
@@ 3.1. Synthetic Data Generation
Since ground-truth causal relationships are typically unavailable in empirical settings, we constructed a synthetic dataset to rigorously assess the accuracy of CCM in both causal discovery and feedback loop identification. The dataset comprises eight time series variables, denoted as $S_{1},\ldots,S_{8}$, each containing 100,000 observations and governed by a lag structure of up to order three. To enable systematic evaluation of feedback detection, we explicitly incorporated bidirectional causal dependencies into the generative process, thereby providing controlled conditions under which the model’s capacity to recover feedback mechanisms can be validated.
@@
Figure 1. True causal graph of the synthetic dataset.
[…]
@@ 3.2. Real-World Dataset and Preprocessing
The real-world dataset consists of 15,584 spatio-temporal records covering Jan 1979–Jun 2021 across the Arctic. Sea ice concentration values (Nimbus-7 SSMR and DMSP SSM/I-SSMIS passive sensors) were combined with 10 atmospheric variables from the ERA-5 reanalysis (Table 1).
@@ 3.3. Benchmark Time Series Causal Discovery Methods
VarLiNGAM. is a Linear Non-Gaussian Acyclic Model for vector autoregressions (VAR), extending LiNGAM to time series. It assumes each variable is a linear function of contemporaneous causes and its own (and others’) past lags, with non-Gaussian, independent errors. Causal discovery separates contemporaneous effects (a DAG) from lagged effects, leveraging ICA-like independence and acyclicity constraints. Estimation fits a VAR, extracts residuals, then applies LiNGAM to residual contemporaneous relations to infer the causal ordering (Jiao et al., 2024).
Proposed CCM Approach We employed the rEDM package (Van Nes et al., 2015), following the workflow in Figure 2. Key steps include: (i) selecting a target variable and computing embeddings $E$ using the simplex function, (ii) choosing optimal $E$ based on forecast skill $\rho$, (iii) applying CCM to detect causal influence and feedback among pairs of variables.
Figure 2. CCM workflow for feedback detection.
[…]
4. Experiments and Results
@@ 4.2.1. Granger Causality
Figure 3. Granger causality test on synthetic data.
[…]
@@ 4.2.2. PCMCI
Figure 4. PCMCI causal graph on synthetic data.
[…]
@@
4.2.3. VarLiNGAM
VarLiNGAM, which integrates LiNGAM with vector autoregressive models, reported significant causal effects up to lag 2. Table 2 summarizes detected links. While it identified several strong causal relations, weaker nonlinear edges were missed. Overall, the model was consistent with the synthetic design but biased toward linear structures.
Table 2. VarLiNGAM statistical test on synthetic data
Source Causal Relationship
S2(t) $\leftarrow$ S1(t-1) (b 0) (100.0%)
S4(t) $\leftarrow$ S1(t-1) (b 0) (67.0%)
S4(t) $\leftarrow$ S3(t) (b 0) (43.0%)
S3(t) $\leftarrow$ S1(t) ($b<0$) (21.0%)
S1(t) $\leftarrow$ S1(t-2) (b 0) (18.0%)
S4(t) $\leftarrow$ S5(t) (b 0) (16.0%)
S3(t) $\leftarrow$ S1(t-1) (b 0) (15.0%)
S1(t) $\leftarrow$ S3(t) ($b<0$) (10.0%)
S1(t) $\leftarrow$ S5(t) (b 0) (3.0%)
S1(t) $\leftarrow$ S1(t-1) (b 0) (2.0%)
S3(t) $\leftarrow$ S5(t) ($b<0$) (2.0%)
S2(t) $\leftarrow$ S1(t) (b 0) (1.0%)
S4(t) $\leftarrow$ S7(t) ($b<0$) (1.0%)
S6(t) $\leftarrow$ S1(t-1) (b 0) (1.0%)

### v5

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; November 03–06, 2025; Minneapolis, MN, USAisbn: 978-1-4503-XXXX-X/2025/12journalyear: 2025copyright: ccconference: The 1st ACM SIGSPATIAL International Workshop on Polar Data Science; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 1st ACM SIGSPATIAL International Workshop on Polar Data Science (PoIDS ’25), November 3–6, 2025, Minneapolis, MN, USAdoi: 10.1145/3764922.3771172isbn: 979-8-4007-2185-4/2025/11ccs: Applied computing Earth and atmospheric sciences
@@ 1. Introduction
The areal extent, concentration, and thickness of sea ice in the Arctic ocean and adjacent seas have dramatically decreased over recent decades (Stroeve et al., 2012). This decline has significantly increased heat flux from the ocean to the atmosphere in autumn and early winter, leading to locally increased air temperatures, moisture, and cloud cover, while reducing static stability in the lower troposphere (Stroeve et al., 2012). Studies suggest that with advancing global warming, cold winters in mid-latitude continents may become uncommon in the second half of the twenty-first century (Blackport et al., 2019). Additionally, recent findings have linked sea ice decline to summer precipitation patterns in Europe, the Mediterranean, and East Asia (Vihma, 2014; Ali et al., 2022). The reduction of Arctic ice sheets is a prominent indicator of warming, evidenced by a nearly 50% decrease in ice extent since 1979 (Stammerjohn et al., 2012). As the Arctic sea ice cover diminishes, the annual mean 2-m air temperature has increased across nearly all weather stations north of 60N, particularly in coastal and archipelago areas surrounding the Arctic Ocean (Polyakov et al., 2012).
@@
The effects of sea ice decline can be categorized into local and remote effects depending on the season. Local effects occur in regions that have lost sea ice cover while remote effects impact areas that have historically been ice-free. The expected direct local effects include increased turbulent fluxes of sensible and latent heat from the ocean to the atmosphere, as well as enhanced longwave radiation emitted by the sea surface. Understanding the remote effects is more complex, as it involves discerning how mid-latitude circulation responds to Arctic changes in the absence of other influencing factors. Interactions among atmospheric time series variables are inherently complex and nonlinear due to the intricate and often unpredictable dependencies that exist between them. In the atmosphere, factors such as temperature, humidity, sea ice extent, and radiation do not evolve in isolation; rather, they interact through multifaceted pathways. For example, an increase in temperature can accelerate sea ice melt, which reduces surface albedo and subsequently amplifies local and regional warming, thereby altering weather patterns. The abundance of feedback loops and deterministic components in atmospheric processes causes information about these variables to become inseparably intertwined. Identifying causal relationships is therefore essential to disentangle these interactions, improve predictive understanding, and design effective interventions. These interactions are fundamentally nonlinear because small perturbations in one variable can trigger disproportionately large responses in others, rendering the system highly sensitive to initial conditions. Moreover, feedback mechanisms where the output of a process feeds back into its own input further amplify complexity and nonlinearity. In this context, uncovering causal relationships and feedback loops is critical for advancing climate science, enabling more accurate forecasts, and informing targeted mitigation and adaptation strategies. Recent work applying Convergent Cross Mapping (CCM) to Arctic datasets has demonstrated its ability to uncover nonlinear feedbacks between sea ice, atmospheric moisture, and radiation processes, validating its use in polar climate research (Su and Wong, 2023; Gao et al., 2023). While controlled randomized experiments are ideal for efficient causal discovery, they are often unethical, costly, or technically unfeasible (Hoyer et al., 2008). As a result, research has focused on causal discovery methods that infer relationships from uncontrolled data (i.e observations collected from natural or real-world processes with many unknown and unmeasured influencing factors). Although stochastic causal models such as Granger causality, PCMCI and VarLiNGAM (they explicitly model causal relationships as probabilistic dependencies driven by random processes) are frequently employed due to their simplicity, they often fail to accurately capture complex, non-linear causal relationships because they are fundamentally constrained by their assumptions of linearity, stationarity, and acyclicity whereas atmospheric and climate dynamics are nonlinear: feedbacks (where processes amplify or dampen their own effects), thresholds (points beyond which behavior qualitatively shifts), or bifurcations (where the system’s regime structure changes abruptly) (Rial et al., 2004), nonstationary (climate change alters statistical properties over time), cyclic and feedback-driven (e.g., ice–albedo, ocean–atmosphere coupling) and finally high-dimensional with hidden confounders. In nonlinear systems every coupled variable carry information about others. This implies that variables cannot be fully removed from the system for analysis, and this further violates the linearity assumption made by stochastic causal models.
@@ 2.1. Background
Detecting causal drivers and regime shifts in Arctic climate systems presents a critical challenge due to nonlinear feedbacks, high-dimensional couplings, and non-separable spatiotemporal dynamics. Sea ice extent, atmospheric circulation, and oceanic heat fluxes interact through bidirectional feedbacks, complicating conventional stochastic causal discovery methods. Convergent Cross Mapping (CCM) addresses these limitations by reconstructing state-space manifolds from observed time series and assessing whether the historical states of one variable can reliably estimate another, enabling detection of weak and asymmetric causal links without relying on predictive models. Concurrently, advances in deep unsupervised spatiotemporal representation learning (Nji et al., 2024b, a; Faruque et al., 2023) demonstrate the ability to uncover latent temporal regimes, structural shifts, and evolving patterns in high-dimensional climate data. While representation learning identifies emergent regime structures, CCM provides mechanistic insight by quantifying the directional couplings that drive these patterns, revealing feedback loops and causal pathways underlying Arctic amplification. Natural systems, such as atmospheric processes, are inherently complex and nonlinear. Linear correlation-based approaches are often inadequate for these systems because correlations can arise without true causal relationships, while genuine causal influences may exist even in the absence of measurable correlation. For causality detection, if $\sigma^{2}(X|U)<\sigma^{2}(X|U-\overline{Y})$, we say that $Y$ is causing $X$, denoted by $Y_{t}\Rightarrow X_{t}$. This means $Y_{t}$ causes $X_{t}$ if we can predict $X_{t}$ better using all available information than without $Y_{t}$.
@@
CCM leverages Takens’ Theorem (Takens, 2006), which states that the hidden dynamics of a complex system can be reconstructed from a single observed time series using time-delayed copies of that variable. This process, known as State-Space Reconstruction, unfolds the system’s trajectory into a multidimensional space where each dimension represents the system’s state at different time lags. Within this reconstructed space, the system evolves along an Attractor Manifold—a geometric surface that captures the stable, recurring patterns toward which the system naturally tends over time. CCM builds upon this principle by constructing “shadow manifolds” from observed variables that preserve the same topological structure as the original system. The embedding dimension $E$ determines how many delayed coordinates are required to fully capture the system’s degrees of freedom, while the time delay $\tau$ controls how separated these snapshots are in time. Choosing appropriate $E$ and $\tau$ ensures that the manifold is unfolded correctly—too small and it collapses, too large and it becomes noisy. Once reconstructed, cross mapping exploits the fact that if variable $X$ causally influences $Y$, then traces of $X$’s dynamics are embedded in $Y$’s manifold, allowing $X$ to be predicted from $Y$. Unlike traditional methods such as Granger causality, which focus on predictive improvement, CCM identifies direct state-space dependence, making it robust for nonlinear and noisy systems. Applications in climate science confirm that CCM successfully identifies causal feedbacks between sea-ice extent, temperature, and radiative processes (Su and Wong, 2023; Gao et al., 2023), offering a powerful framework for disentangling drivers of Arctic sea-ice decline and associated regime shifts.
@@ 2.2. Related Work
Convergent Cross Mapping is designed to identify and quantify causalities in systems whose variables are not separable. Clark et al. proposed Multispatial CCM which combine the existing techniques of CCM and dewdrop regression to build a novel test of causal relations that leverages spatial replication to detect causal relationships between processes (Clark et al., 2015). While the ability of their proposed model was tested on simulated and real-world ecological data, their performance on real world spatiotemporal data is yet unexplored. Brouwer et al. (De Brouwer et al., 2020), proposed Latent CCM which uses reconstruction between latent processes of dynamical systems to infer causality between short and sporadic time series. Although their proposed seemed to work well on neural activity data, its application to climate data has not yet been explored. Ye et. al. (Ye et al., 2015), demonstrate the ability of CCM to identify different time-delayed interactions, distinguish between synchrony induced by strong unidirectional-forcing and true bidirectional causality, and resolve transitive causal chains. Their application is limited to model simulations, a laboratory predator-prey experiment, temperature and greenhouse gas reconstructions from the Vostok ice core, and longterm ecological time series collected in the Southern California Bight with no known application to complex multidimensional atmospheric data. Wang et. al. applied CCM to explore the causality between soil moisture and precipitation over low- and mid- latitude regions (Wang et al., 2018). Similarly, recent Arctic-focused studies demonstrate CCM’s strength in resolving feedback loops between sea ice extent and atmospheric drivers such as humidity and radiation (Gao et al., 2023; Ye et al., 2015).
@@ 3.1. Synthetic Data Generation
Our goal is to detect both linear and non-linear causal relationships, including weak and bidirectional feedback loops, among time series atmospheric variables. To achieve this, we evaluate a set of causal discovery methods without bias: stochastic models (Granger causality (Barnett et al., 2009), PCMCI (Runge, 2020), VarLiNGAM (Shimizu et al., 2006)) and the nonlinear Convergent Cross Mapping (CCM), applied to both synthetic and real-world Arctic climate data. Since ground-truth causal relationships are typically unavailable in empirical settings, we constructed a synthetic dataset to rigorously assess the accuracy of CCM in both causal discovery and feedback loop identification. The dataset comprises eight time series variables, denoted as $S_{1},\ldots,S_{8}$, each containing 100,000 observations and governed by a lag structure of up to order three. To enable systematic evaluation of feedback detection, we explicitly incorporated bidirectional causal dependencies into the generative process, thereby providing controlled conditions under which the model’s capacity to recover feedback mechanisms can be validated.
@@
Figure 1. True causal graph of the synthetic dataset.:
@@ 3.2. Real-World Dataset and Preprocessing
Table 1 consists of real-world Copernicus Arctic Regional ReAnalysis (CARRA) dataset and provides high-resolution spatiotemporal climate variables over the Arctic domain. The temporal resolution is hourly and aggregated to daily averages in this dataset, covering the full calendar year of 2019 (January 1–December 31), allowing fine-grained analysis of day-to-day and seasonal variability. Spatially, the data are gridded on a 0.03° × 0.03° ( 3 km × 3 km) horizontal resolution, representing a detailed depiction of mesoscale Arctic processes. Each grid cell records atmospheric and surface parameters—such as temperature, snow depth, wind components, and radiation fluxes—capturing both temporal evolution and spatial heterogeneity across Greenland and the Arctic region. This high-resolution spatiotemporal coverage enables accurate regional climate and snowmelt pattern analyses.
@@ 3.3. Benchmark Time Series Causal Discovery Methods
VarLiNGAM is a Linear Non-Gaussian Acyclic Model for vector autoregressions (VAR), extending LiNGAM to time series. It assumes each variable is a linear function of contemporaneous causes and its own (and others’) past lags, with non-Gaussian, independent errors. Causal discovery separates contemporaneous effects (a DAG) from lagged effects, leveraging ICA-like independence and acyclicity constraints. Estimation fits a VAR, extracts residuals, then applies LiNGAM to residual contemporaneous relations to infer the causal ordering (Jiao et al., 2024).
@@ 3.4. Hardware Setup
Proposed CCM Approach.
Figure 2. CCM workflow for feedback detection.:
We employed the rEDM package (Van Nes et al., 2015), following the workflow in Figure 2. Key steps include: (i) selecting a target variable and computing embeddings $E$ using the simplex forecasting function (even for a chaotic time series, future values can be predicted from the behaviour of similar past values) (ii) choosing optimal $E$ based on forecast skill $\rho$, (iii) applying CCM to detect causal influence and feedback among pairs of variables.
4. Experiments and Results
@@ 4.1.1. Granger Causality
Figure 3. Learned causal graph from synthetic data using Granger causality method.:
@@ 4.1.2. PCMCI
Figure 4. Learned causal graph from synthetic data using PCMCI.:
@@
4.1.3. VarLiNGAM.
It integrates LiNGAM with vector autoregressive models, reported significant causal effects up to lag 2. Figure 5 is the learned causal graph from the synthetic data. Nodes are represented only with their variable names with lags collapsed for visualization purposes. We applied a p-value lower threshold of 0.05 and edges below this threshold were dropped. Overall, the model was consistent with the synthetic design but biased toward linear structures.
Figure 5. Learned causal graph from synthetic data using VarLiNGAM.: